# LLM-SQL-Solver: Can LLMs Determine SQL Equivalence?

### v3

@@ 5 Discussion and Limitation
Concurrent with our research, the chain-of-table approach Wang et al. (2024) independently explores a similar concept, allowing LLMs to form a reasoning path through table modifications to improve question answering from tabular data, akin to our Miniature & Mull prompt. However, while our technique focuses on determining query semantic equivalence through table creation and modification, their goal is to deepen the understanding of the tabular data itself.
There are some limitations or areas of improvement to this work. First, we only evaluated the capabilities of the Large Language Models of OpenAI. There are other alternative models such as Llama2 7B Touvron et al. (2023), Vicuna Chiang et al. (2023), Alpaca Taori et al. (2023a) which also showcase strong understanding and reasoning for logical tasks Chen et al. (2021); Yue et al. (2023); Yan et al. (2023b). Nevertheless, we believe that GPT 3.5 turbo and GPT 4 turbo exemplify the current status of the GPT model family’s capabilities. In addition, our evaluations on Spider Benchmark were conducted over randomly sampled questions, in which we have to manually label the equivalence of the two different notions (semantic and relaxed). Since there are nearly ten thousand pairs of SQL queries, manually checking all of them are arguably infeasible. In fact, even checking a pair of SQL queries for semantic view takes around 10 minutes for experts in SQL. Although the questions are sampled, we observe a consistent trends between LLMs’ judgements and execution matches (such as GPT 4 turbo’s responses are consistently better aligned with human preferences).
@@ 6 Conclusion
In this paper, we studied and answered two main questions: i) Can current LLMs be used to help engineers in checking semantic equivalences between SQL queries and ii) should we use LLMs to judge the quality of generated SQL queries for text-to-SQL benchmarks. To the best of our knowledge, this is the first work studying the ability of LLMs in understanding SQL queries and determining SQL equivalences. Through extensive evaluations and manually examining the quality of the answers based on the alignment with human preferences, we find current LLMs (especially GPT 4 turbo), using our proposed prompt techniques, can assists engineers in determining semantic equivalent queries, but may yield incorrect responses. On the other hand, when comparing the generated SQL query with the gold SQL query with our relaxed notion of equivalence, LLMs’ responses constantly achieve high alignments with human preferences. Moreover, our proposed LLM-SQL-Solver avoids accessing stored user data, and hence no sophisticated databases need to be constructed.

### v4

@@ 5 Discussion and Future Works
In this section, we would like to highlight the broader implications and potential integration of our proposed LLM-SQL-Solver. Moreover, Concurrent with our research, the chain-of-table approach Wang et al. (2024) independently explores a similar concept, allowing LLMs to form a reasoning path through table modifications to improve question answering from tabular data, akin to our Miniature & Mull prompt. However, while our technique focuses on determining query semantic equivalence through table creation and modification, their goal is to deepen the understanding of the tabular data itself.
5.1 Applications of LLM-SQL-Solver
First, LLM-SQL-Solver (semantic notion) can be directly used to find system bugs (e.g., COUNT BUG) and help engineers validate their SQL rewriting code bases. Second, since query optimizer relies on heuristic rules to find equivalent SQL forms, it may not explore all rewriting possibilities and hence, the rewritten query may lead to sub-optimal performance. We believe that during the SQL rewrite phase of the query optimization engine, the LLM-SQL-Solver can be directly used as a tool to help checking semantic equivalence, ensuring bugs (e.g., COUNT bug Ganski and Wong (1987)) do not occur and also explore more SQL rewrite possibilities which may lead to better query performance.
For text-to-SQL tasks, we believe that when comparing the gold SQL query (the correct answer) with the generated SQL query, the current most popular metric, execution match, does not reflects the human perspective and fails to capture a user’s opinion on the generated SQL. For instance, a common reasoning for failures in execution match is that the generated SQL selects fewer/extra columns than the gold SQL. In fact, a recent work has explicitly added "do not select extra columns" in its prompt to improve the execution accuracy Dong et al. (2023). However, from a user’s perspective providing extra columns may not reduce the quality of the answer. Also, there are many other similar reasons that can lead to false execution matches but not necessarily indicate a lower quality of the generated SQL query (such as the column order in selection, case sensitivity, duplicate entries, and columns in different tables share the similar contents). Moreover, execution matching fails to correctly identify logically different queries when the underlying database does not contain any counter examples, even when hundreds of human hours were consumed to create the benchmark Yu et al. (2018). Lastly, a fundamental limitation of execution accuracy is its need to query the underlying database. This is particularly pertinent as the text-to-SQL service provider usually does not have ownership of the user’s database and the user’s database may also contain sensitive information. As a result, it would be desirable to have a SQL solver determining the relax notion of equivalence without constructing sophisticated databases (save costs) and without accessing any stored data (ensure data privacy). Based on our extensive experiment evaluations, we believe that LLM-SQL-Solver (with pragmatic notion) can be used as a better metric for text-to-SQL benchmarks.
5.2 Future Plans
There are areas of improvement to this work. First, we find LLMs may hallucinate in executing the queries over simple databases. By equipping LLMs with tools such as Python pandas library, LLMs can reply on these tool for observing the execution results of SQL queries Zhuang et al. (2023). By reducing the hullicination, we believe that our proposed Miniature & Mull prompt will have an increased accuracy in determining SQL semantic equivalence.
In addition, our evaluations of pragmatic equivalence were conducted on a set of 70 questions, requiring experts to manually label the equivalence of each pair of SQL queries. This process is time-consuming, as even experienced SQL experts need approximately 10 minutes to thoroughly analyze and determine the equivalence of a single query pair. Despite the relatively small dataset size, we observed consistent trends in the alignment between LLM judgments and execution matches. Notably, GPT-4o’s responses consistently demonstrated better alignment with human preferences compared to other models, showcasing its ability to better capture nuanced distinctions in query intent. Expanding the dataset would allow for a more robust analysis and further validation of these trends, providing deeper insights into the strengths and weaknesses of current models.
Moreover, our analysis revealed that pragmatic equivalence is often highly subjective, depending on the context and the user’s interpretation of query intent. This subjectivity underscores the need for designing customizable evaluation metrics tailored to individual users. Such personalized metrics could adapt to specific user needs and preferences, ultimately improving the quality and relevance of text-to-SQL systems for diverse applications. Future research in this direction could significantly enhance the usability and practicality of these systems, enabling them to cater to a broader range of use cases and user requirements.
@@ 6 Conclusion
In this paper, we studied and answered two main questions: i) Can current LLMs be used to help engineers in checking semantic equivalences between SQL queries and ii) should we use LLMs to judge the quality of generated SQL queries for text-to-SQL benchmarks. To the best of our knowledge, this is the first work studying the ability of LLMs in understanding SQL queries and determining SQL equivalences. Through extensive evaluations and manually examining the quality of the answers based on the alignment with human preferences, we find current LLMs (especially GPT-4o), using our proposed prompt techniques, can assists engineers in determining semantic equivalent queries, but may yield incorrect responses. On the other hand, when comparing the generated SQL query with the gold SQL query with pragmatic equivalence notion, LLMs’ responses constantly achieve high alignments with human preferences. Moreover, using our proposed LLM-SQL-Solver as the evaluation metric for text-to-SQL translation ensures data privacy by eliminating the need to access stored user data, and it reduces costs by removing the requirement to construct sophisticated databases, making the evaluation process more efficient and scalable.
Acknowledgment
We thank Tom Shangguan for his assistance in organizing and managing the data files used in this research.